# A Momentum Balance Correction to the Non-Conservative One-Fluid Formulation in Boiling Flows using Volume-of-Fluid

### v1

@@ 3.4 Formal Analysis of the Modified Predictor-Projection Method
where I, G and D are the discrete identity, gradient and divergence operators, respectively. Further, $\mathbf{R}$ corresponds to all the discretized terms on the RHS of the momentum equation, Eq. (15), plus the convective term. The various body forces are included here. That is, for the purpose of this analysis, no gradient splitting operator is implemented nor the body forces are moved to the projector step. Moreover, the temporal scheme for $\mathbf{R}$ is left unspecified and could be obtained from, e.g., an explicit forward Euler or from the AB scheme. The volumetric expansion at the interface is given by $\text{M}^{n+1}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}|^{n+1}$. Note that the predictor-projection steps can be obtained by performing the LU decomposition of $\mathbf{A}\mathbf{x}=\mathbf{b}$ [52, 53], i.e., a Poisson-type equation emerges naturally.
@@
To improve the numerical solution of Eq. (28), $\mathbf{u^{n}}$ is regularized by introducing a Stefan flow shift before proceeding to the FSM, as discussed in Section 3.3. Such a shift removes the dependence of pressure with the previous time step by introducing an intermediate velocity field, $\mathbf{u}^{*}$, which corrects the velocity field $\mathbf{u}^{n}$ to satisfy the volumetric expansion at the next time step $\text{M}^{n+1}$. The shift is defined implicitly by introducing an intermediate pressure jump, $p^{*}$, resulting in an additional linear system of equations given by
@@
$\text{D}\frac{1}{\rho}\text{G}p^{n+1}=\text{D}\mathbf{u^{*}}-\text{M}^{n+1}=\Delta t\text{D}\mathbf{R}+\text{D}\mathbf{u^{*}}-\text{M}^{n+1}$ (33)
@@
Various details of the algorithm have become clear with the previous analysis and are summarized here. First, the flow solver has been split into two predictor-projection or FSM steps rather than one. By doing so, one integrates the velocity first from $\mathbf{u}^{n}$ to $\mathbf{u}^{*}$ purely based on mass conservation constraints due to the moving interface, absorbing the time oscillations in the pressure term $p^{*}$; then, one integrates momentum from $\mathbf{u}^{*}$ to $\mathbf{u}^{n+1}$, as visualized in Figure 3. Note that $\mathbf{R}$ is discretized with $\mathbf{u}^{*}$ as shown in Eq. (24). Similar approaches with intermediate velocities are used to achieve other goals, such as momentum-consistent advection in [17]. Second, by using the first FSM step, one eliminates the divergence constraint in $\Omega_{\Gamma}$ and obtains an oscillation-free pressure field $p^{n+1}$. Third and last, the recasting of Eq. (32) shows that the second FSM step is indeed solving for a new pressure $p^{n+1}=p^{n+1}_{\text{FSM}}-p^{*}$ that differs from the pressure field that one would obtain from directly solving the system given by Eq. (28) by $p^{*}$. In other words, $p^{*}$ is an additional pressure added to the physical pressure field resulting from an inconsistent temporal integration between the Navier-Stokes equations and the interface advection when $\dot{m}^{\prime}\neq 0$.
@@
A formal analysis of the impact on the temporal accuracy of using $\mathbf{u}^{*}$ instead of $\mathbf{u}^{n}$ in the integration of $\mathbf{R}$ is out the scope of this paper. However, since changes in the velocity field caused by the Stefan flow shift are expected to be mainly confined in the vicinity of the interface $\Omega_{\Gamma}$, where jump conditions are dominant over inertial terms, this is assumed to be negligible given a small enough time interval.
@@ 3.5 One-dimensional Discrete Analysis of Momentum Balance Corrections
A one-dimensional discrete analysis of the momentum balance corrections is provided in this section to visualize the modifications introduced in the flow solver discussed in Sections 3.3 and 3.4. An evaporating liquid film sitting on a fixed wall is analyzed where the liquid is at rest and the flat surface does not deform. Similar to the Stefan problem but without coupling to the energy equation [13, 17, 35], this configuration results in the one-dimensional evolution of an evaporating flow. The analytical solution is straightforward, where the velocities in the liquid and gas phases are constant and given by $u_{l}=0$ and $u_{g}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Thus, the viscous term is identically zero. Moreover, phase change causes a pressure jump across the interface given by $\Delta p=(\dot{m}^{\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$. The discretization of the one-dimensional domain is shown in Figure 4.
@@
From this setup one can directly show the effects of the two modifications proposed to the flow solver, namely the intermediate step to calculate $\mathbf{u}^{*}$ and the addition of $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$. Combining Eqs. (24) and (26) in $\Omega_{\Gamma}$ where the forward Euler scheme is used for the temporal integration, the pressure gradient across a velocity node, e.g., $i$+$\tfrac{1}{2}$, is given by
$\frac{dp^{n+1}}{dx}=\frac{\rho_{0}}{\Delta t^{n+1}}(u^{*}-u^{n+1})-\rho_{0}u^{*}\frac{du^{*}}{dx}+\bigg{(}1-\frac{\rho_{0}}{\rho^{n+1}}\bigg{)}\bigg{(}\frac{d\hat{p}}{dx}-\hat{f}_{\dot{m}^{\prime},x}\bigg{)}+f_{\dot{m}^{\prime},x}^{n+1}+\frac{\rho_{0}}{\rho^{n+1}}f_{NC,x}^{n+1}$ (34)
The pressure field obtained from predictor-projection methods is essentially a mathematical operation to enforce the divergence-free condition of the velocity field in incompressible flows. As long as $\mathbf{u}^{**}$ is well posed, the pressure is physically meaningful. However, this is not the case with the one-fluid formulation introduced in Section 2.3 across $\Omega_{\Gamma}$ if $\dot{m}\neq 0$. Fundamentally, the PPE is used to overwrite the velocity field in $\Omega_{\Gamma}$ to satisfy Eq. (10) regardless of $\mathbf{u}^{**}$. Thus, the correct velocity field may be obtained but with an ill-defined momentum balance. The introduction of $\mathbf{u}^{*}$, $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime}}$ aim to solve this issue by cancellation and correction. Similar to the modeling of surface tension in [50], the discretization of $f_{\dot{m}^{\prime},x}^{n+1}$ and $\hat{f}_{\dot{m}^{\prime},x}$ consistent with the pressure gradient ensures that the exact pressure jump due to phase change is recovered across $\Omega_{\Gamma}$ if no other jumps are introduced numerically.
@@
where $u_{i-\frac{1}{2}}^{*}=u_{l}=0$ and $u_{i+\frac{5}{2}}^{*}=u_{g}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Further, from $\nabla\cdot\mathbf{u}^{*}$, one obtains
$\begin{split}&u_{i+\frac{1}{2}}^{*}=\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i}\Delta x\\ &u_{i+\frac{3}{2}}^{*}=u_{i+\frac{1}{2}}^{*}+\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i+1}\Delta x\\ &u_{i+\frac{5}{2}}^{*}=u_{i+\frac{3}{2}}^{*}+\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i+2}\Delta x\end{split}$ (36)
@@
$u^{*}\frac{du^{*}}{dx}\approx\begin{cases}u_{i+\frac{1}{2}}^{*}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\frac{1}{2}(\delta_{\Gamma,i}+\delta_{\Gamma,i+1})=-u_{i+\frac{1}{2}}^{*}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\bigg{(}\frac{C_{i+\frac{3}{2}}-C_{i-\frac{1}{2}}}{2\Delta x}\bigg{)}&\text{if $i+\frac{1}{2}$}\\ u_{i+\frac{3}{2}}^{*}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\frac{1}{2}(\delta_{\Gamma,i+1}+\delta_{\Gamma,i+2})=-u_{i+\frac{3}{2}}^{*}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\bigg{(}\frac{C_{i+\frac{5}{2}}-C_{i+\frac{1}{2}}}{2\Delta x}\bigg{)}&\text{if $i+\frac{3}{2}$}\\ \end{cases}$ (37)
@@
The adjustment of the Stefan flow with the first predictor-projection step effectively eliminates the pressure jump introduced by the time derivative in Eqs. (11) or (15), i.e., the first term on the LHS of Eq. (12). Without correction, $u^{*}=u^{n}$ and the temporal term in Eq. (34) becomes $\frac{\rho_{0}}{\Delta t^{n+1}}(u^{n}-u^{n+1})$. In, e.g., node $i$+$\tfrac{1}{2}$, it is given by
$\frac{\rho_{0}}{\Delta t^{n+1}}(u^{n}-u^{n+1})=-\frac{\rho_{0}}{\Delta t^{n+1}}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})(\delta_{\Gamma,i}^{n+1}-\delta_{\Gamma,i}^{n})\Delta x$ (38)
Note that if $\rho(\mathbf{u}_{L}-\mathbf{u}_{G})(\mathbf{u}_{\Gamma}\cdot\mathbf{n}_{\Gamma})$ from Eq. (12) is rewritten as a localized body force similar to $\mathbf{f}_{NC}$, it would be given in node $i$+$\tfrac{1}{2}$ as $-\rho_{0}\dot{m}^{\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})u_{\Gamma,i+\frac{1}{2}}^{n+1}\delta_{\Gamma,i+\frac{1}{2}}^{n+1}$. Thus, one can see that the effect of the interface shift over time on the momentum balance of the non-conservative formulation, driven by $u_{\Gamma,i+\frac{1}{2}}^{n+1}\delta_{\Gamma,i+\frac{1}{2}}^{n+1}$, is approximated in the predictor-projection method with $(\delta_{\Gamma,i}^{n+1}-\delta_{\Gamma,i}^{n})\frac{\Delta x}{\Delta t^{n+1}}$. That is, the interface velocity is approximated by the shift in $\delta_{\Gamma}$ from $t^{n}$ to $t^{n+1}$.
@@ 4 Validation with Fixed Mass Flux
The numerical approach and proposed momentum balance corrections are validated against fabricated solutions of evaporating flows with constant $\dot{m}^{\prime}$. This effectively decouples the flow solver from the energy equation. Thus, a better assessment of the handling of phase change in the advection of the interface and the Navier-Stokes equations can be made. Moreover, gravity is neglected.
@@
Note that $\dot{m}^{\prime}$ values are pushed to what could be considered upper limits in evaporative flows. Since the modifications to what is considered a standard two-phase flow solver disappear as $\dot{m}^{\prime}\rightarrow 0$, it is out of scope of this work to show results approaching non-evaporative conditions. Further, in this section the CFL constraints on the time step are driven by the interface evaporation. As such, a restrictive CFL condition on the interface shift limits the maximum shift per time step to $(\Delta\text{d}_{\text{max}}/\Delta x)=0.001$ to minimize the geometrical errors introduced by the respective term in Eq. (9) and focus on the behavior of the additional body forces $\mathbf{f}_{\dot{m}^{\prime}}$ and $\mathbf{f}_{NC}$.

### v2

@@ 3.4 Formal Analysis of the Modified Predictor-Projection Method
where I, G and D are the discrete identity, gradient and divergence operators, respectively. Further, $\mathbf{R}$ corresponds to all the discretized terms on the RHS of the momentum equation, Eq. (15), plus the convective term. The various body forces are included here. That is, for the purpose of this analysis, no gradient splitting operator is implemented nor the body forces are moved to the projector step. Moreover, the temporal scheme for $\mathbf{R}$ is left unspecified and could be obtained from, e.g., an explicit forward Euler or from the AB scheme. The volumetric expansion at the interface is given by $\text{M}^{n+1}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma}|^{n+1}$. Note that the predictor-projection steps can be obtained by performing the LU decomposition of $\mathbf{A}\mathbf{x}=\mathbf{b}$ [54, 55], i.e., a Poisson-type equation emerges naturally.
@@
To improve the numerical solution of Eq. (28), $\mathbf{u^{n}}$ is regularized by introducing a Stefan flow shift before proceeding to the FSM, as discussed in Section 3.3. Such a shift removes the dependence of pressure with the Stefan flow of the previous time step by introducing an intermediate velocity field, $\mathbf{u}^{*}$, which corrects the velocity field $\mathbf{u}^{n}$ to satisfy the volumetric expansion at the next time step $\text{M}^{n+1}$. The shift is defined implicitly by introducing an intermediate pressure, $p^{*}$, resulting in an additional linear system of equations given by
@@
$\text{D}\frac{1}{\rho}\text{G}p^{n+1}=\text{D}\mathbf{u^{**}}-\text{M}^{n+1}=\Delta t\text{D}\mathbf{R}+\text{D}\mathbf{u^{*}}-\text{M}^{n+1}$ (33)
@@
Various details of the algorithm have become clear with the previous analysis and are summarized here. First, the flow solver has been split into two predictor-projection or FSM steps rather than one. By doing so, one integrates the velocity first from $\mathbf{u}^{n}$ to $\mathbf{u}^{*}$ purely based on mass conservation constraints due to the moving interface, absorbing the time oscillations in the pressure term $p^{*}$; then, one integrates momentum from $\mathbf{u}^{*}$ to $\mathbf{u}^{n+1}$, as visualized in Figure 3. Second, by using the first FSM step, one eliminates the divergence constraint in $\Omega_{\Gamma}$ and obtains a pressure field $p^{n+1}$ free of oscillations caused by a time-dependent stiff source term, resembling the implementation of the FSM in an incompressible single-phase problem. Third and last, the recasting of Eq. (32) shows that the second FSM step is indeed solving for a new pressure $p^{n+1}=p^{n+1}_{\text{FSM}}-p^{*}$ that differs from the pressure field that one would obtain from directly solving the system given by Eq. (28). In other words, the numerical pressure $p^{n+1}_{\text{FSM}}$ obtained with the traditional formulation of the FSM includes an additional pressure $p^{*}$ resulting from the temporal change of the Stefan flow, which carries the problematic pressure behavior across $\Omega_{\Gamma}$ caused by the inconsistent temporal integration between the one-fluid Navier-Stokes equations and the interface transport when $\dot{m}^{\prime\prime}\neq 0$.
@@
Some limitations arise with the proposed two-step predictor-projection method. As shown in Eq. (24), $\mathbf{R}$ is discretized with $\mathbf{u}^{*}$ instead of $\mathbf{u}^{n}$. Thus, the impact on the evaluation of $\mathbf{R}$ and how it might affect the temporal integration of momentum must be understood. This is left for future work, but preliminary analysis following the static droplet evaporation case presented in Section 4.2 suggests that the convective and diffusive terms differ by less than 0.03% given the considered $\Delta t$ restrictions. Further, by removing the information carried by $p^{*}$ regarding unsteady effects of the Stefan flow, one effectively imposes that the interface movement with respect to the fluid is a quasi-steady state process. Therefore, the pressure $p^{n+1}$ obtained with the modified FSM is only physical if the quasi-steady assumption holds despite the correct pressure jump and volumetric expansion may be recovered across $\Omega_{\Gamma}$ when solving the Navier-Stokes equations. While this assumption may not be problematic in cases where the flow dynamics are driven by other forces (e.g., gravity), differences become more evident if the flow field is dominated by the evaporation process.
@@ 3.5 One-dimensional Discrete Analysis of Momentum Balance Corrections
A one-dimensional discrete analysis of the momentum balance corrections is provided in this section to visualize the modifications introduced in the flow solver discussed in Sections 3.3 and 3.4. An evaporating liquid film sitting on a fixed wall is analyzed where the liquid is at rest and the flat surface does not deform. Similar to the Stefan problem but without coupling to the energy equation [13, 17, 35], this configuration results in the one-dimensional evolution of an evaporating flow. The analytical solution is straightforward, where the velocities in the liquid and gas phases are constant and given by $u_{l}=0$ and $u_{g}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Thus, the viscous term is identically zero and the quasi-steady assumption of the interface regression with respect to the fluid has no impact beyond $\Omega_{\Gamma}$. Here, phase change causes a pressure jump across the interface given by $\Delta p=(\dot{m}^{\prime\prime})^{2}(\rho_{G}^{-1}-\rho_{L}^{-1})$. The discretization of the one-dimensional domain is shown in Figure 4.
@@
From this setup one can directly show the effects of the two modifications proposed to the flow solver, namely the intermediate step to calculate $\mathbf{u}^{*}$ and the addition of $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$. Combining Eqs. (24) and (26) in $\Omega_{\Gamma}$ where the forward Euler scheme is used for the temporal integration, the pressure gradient across a velocity node, e.g., $i$+$\tfrac{1}{2}$, is given by
$\frac{dp^{n+1}}{dx}=\frac{\rho_{0}}{\Delta t^{n+1}}(u^{*}-u^{n+1})-\rho_{0}u^{*}\frac{du^{*}}{dx}+\bigg{(}1-\frac{\rho_{0}}{\rho^{n+1}}\bigg{)}\bigg{(}\frac{d\hat{p}}{dx}-\hat{f}_{\dot{m}^{\prime\prime},x}\bigg{)}+f_{\dot{m}^{\prime\prime},x}^{n+1}+\frac{\rho_{0}}{\rho^{n+1}}f_{NC,x}^{n+1}$ (34)
The pressure field obtained from predictor-projection methods is essentially a mathematical operation to enforce the divergence-free condition of the velocity field in incompressible flows. As long as $\mathbf{u}^{**}$ is well posed, the pressure is physically meaningful. However, this is not the case with the one-fluid formulation introduced in Section 2.3 across $\Omega_{\Gamma}$ if $\dot{m}^{\prime\prime\prime}\neq 0$. Fundamentally, the PPE is used to overwrite the velocity field in $\Omega_{\Gamma}$ to satisfy Eq. (10) regardless of $\mathbf{u}^{**}$. Thus, the correct velocity field may be obtained but with an ill-defined momentum balance. The introduction of $\mathbf{u}^{*}$, $\mathbf{f}_{NC}$ and $\mathbf{f}_{\dot{m}^{\prime\prime}}$ aim to solve this issue by cancellation and correction. Similar to the modeling of surface tension in [51], the discretization of $f_{\dot{m}^{\prime\prime},x}^{n+1}$ and $\hat{f}_{\dot{m}^{\prime\prime},x}$ consistent with the pressure gradient ensures that the exact pressure jump due to phase change is recovered across $\Omega_{\Gamma}$ if no other jumps are introduced numerically.
@@
where $u_{i-\frac{1}{2}}^{*}=u_{l}=0$ and $u_{i+\frac{5}{2}}^{*}=u_{g}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})$. Further, from $\nabla\cdot\mathbf{u}^{*}$, one obtains
$\begin{split}&u_{i+\frac{1}{2}}^{*}=\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i}\Delta x\\ &u_{i+\frac{3}{2}}^{*}=u_{i+\frac{1}{2}}^{*}+\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i+1}\Delta x\\ &u_{i+\frac{5}{2}}^{*}=u_{i+\frac{3}{2}}^{*}+\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\delta_{\Gamma,i+2}\Delta x\end{split}$ (36)
@@
$u^{*}\frac{du^{*}}{dx}\approx\begin{cases}u_{i+\frac{1}{2}}^{*}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\frac{1}{2}(\delta_{\Gamma,i}+\delta_{\Gamma,i+1})=-u_{i+\frac{1}{2}}^{*}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\bigg{(}\frac{C_{i+\frac{3}{2}}-C_{i-\frac{1}{2}}}{2\Delta x}\bigg{)}&\text{if $i+\frac{1}{2}$}\\ u_{i+\frac{3}{2}}^{*}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\frac{1}{2}(\delta_{\Gamma,i+1}+\delta_{\Gamma,i+2})=-u_{i+\frac{3}{2}}^{*}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})\bigg{(}\frac{C_{i+\frac{5}{2}}-C_{i+\frac{1}{2}}}{2\Delta x}\bigg{)}&\text{if $i+\frac{3}{2}$}\\ \end{cases}$ (37)
@@
The adjustment of the Stefan flow with the first predictor-projection step eliminates the pressure jump introduced by the time derivative in Eqs. (11) or (15), i.e., the first term on the LHS of Eq. (12). Without correction, the temporal term in Eq. (34) causing a pressure jump becomes $\frac{\rho_{0}}{\Delta t^{n+1}}(u^{n}-u^{n+1})$. In, e.g., node $i$+$\tfrac{1}{2}$, it is given by
$\frac{\rho_{0}}{\Delta t^{n+1}}(u^{n}-u^{n+1})=-\frac{\rho_{0}}{\Delta t^{n+1}}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})(\delta_{\Gamma,i}^{n+1}-\delta_{\Gamma,i}^{n})\Delta x$ (38)
Note that if $\rho(\mathbf{u}_{L}-\mathbf{u}_{G})(\mathbf{u}_{\Gamma}\cdot\mathbf{n}_{\Gamma})$ from Eq. (12) is rewritten as a localized body force similar to $\mathbf{f}_{NC}$, it could be approximated at $i$+$\tfrac{1}{2}$ as $-\rho_{0}\dot{m}^{\prime\prime}(\rho_{G}^{-1}-\rho_{L}^{-1})u_{\Gamma,i+\frac{1}{2}}^{n+1}\delta_{\Gamma,i+\frac{1}{2}}^{n+1}$. Thus, the effect of the interface shift over time on the momentum balance of the non-conservative formulation, driven by $u_{\Gamma,i+\frac{1}{2}}^{n+1}\delta_{\Gamma,i+\frac{1}{2}}^{n+1}$, is approximated in the predictor-projection method with $(\delta_{\Gamma,i}^{n+1}-\delta_{\Gamma,i}^{n})\frac{\Delta x}{\Delta t^{n+1}}$. That is, the interface velocity is approximated by the shift in $\delta_{\Gamma}$ from $t^{n}$ to $t^{n+1}$, which is eliminated by considering the intermediate velocity $\mathbf{u}^{*}$ instead.
@@ 4 Validation with Fixed Mass Flux
The numerical approach and proposed momentum balance corrections are validated against fabricated solutions of evaporating flows with constant $\dot{m}^{\prime\prime}$. This effectively decouples the flow solver from the energy equation. Thus, a better assessment of the handling of phase change in the advection of the interface and the Navier-Stokes equations can be made. Moreover, gravity is neglected.
@@
Note that $\dot{m}^{\prime\prime}$ values are pushed to what could be considered upper limits in evaporative flows. Since the modifications to what is considered a standard two-phase flow solver disappear as $\dot{m}^{\prime\prime}\rightarrow 0$, it is out of scope of this section to show results approaching non-evaporative conditions. Further, in this section the CFL constraints on the time step are driven by the interface evaporation. As such, a restrictive CFL condition on the interface shift limits the maximum shift per time step to $(\Delta\text{d}_{\text{max}}/\Delta x)=0.001$ to minimize the geometrical errors introduced by the respective term in Eq. (9) and focus on the behavior of the additional body forces $\mathbf{f}_{\dot{m}^{\prime\prime}}$ and $\mathbf{f}_{NC}$.